# An LLM-based Agent for Reliable Docker Environment Configuration

### v2

@@ 3.1 External environment
3.2 Internal environment
As shown in the Repo2Run workflow of Figure 3, the internal environment consists of a Docker container with a base image $B$ and the code repository to be configured. According to the latest data [^1] https://w3techs.com/technologies/history_details/pl-python/3 from 2025, we select Python 3.10 as our default Docker base image due to its increasing usage and wide application among all Python 3 versions. Additionally, we categorize the executable actions within the internal environment into five types: environment monitoring, dependency installation, bash commands, code editing, and test running.
$\bullet$ Environment monitoring: It serves as the eyes of the LLM agent in the internal environment, allowing it to observe the current state of the environment. These commands usually do not change the state of the environment. For instance, commands like “ls” and “cat” are employed to inspect directories and files. Additionally, commands like “pip list” and “pipdeptree [^2] https://github.com/tox-dev/pipdeptree are utilized to get the installed version of third-party libraries in the current environment. For example, in the fourth part of Event Stream in Figure 3, the LLM agent calls the “find” command to search for requirements.txt within the environment to determine which dependencies need to be installed.
$\bullet$ Dependency installation: It is primarily used to install the third-party libraries necessary for running tests, including libraries managed by both “pip” and “apt-get”. To avoid dependency conflicts that may arise when downloading multiple packages, we implement a dependency management system in the external environment. Initially, the third-party libraries to be downloaded are added to a waiting list. If there are different constraints for the same library (e.g., “A>=>=> =1.0” and “A$<$1.0”), they will be moved to a conflict list, and the LLM agent will determine the appropriate version to download.
We design a set of tools for the LLM agent to manage the installations. For instance, as shown in Figure 3, the action “waitinglist addfile” adds all elements from requirements.txt to the waiting list, while executing the “download” command, third-party libraries are taken one by one from the waiting list and installed in the internal environment. The installation commands (e.g., “pip install” and “apt-get install”) are executed. If the installation succeeds, the state of the environment is updated; if it fails, a rollback is performed. We show all our designed tools in Appendix C.
$\bullet$ Test running: It detects whether the current environment is successfully configured by running “pytest” unit tests. If the state ($S_{f}$) of the current environment passes the tests, it indicates that the LLM agent has successfully configured the environment within the Docker container, and the internal environment’s process concludes. Conversely, if some tests fail, the error information is communicated back to the LLM agent, allowing it to perform further configurations and adjustments based on the errors.
$\bullet$ Code editing: It allows the LLM agent to modify the code within the internal environment, including codes inside and outside the repository. Besides, to prevent the LLM agent from spuriously running the tests by directly modifying or deleting the test files, we do not permit the LLM agent to modify or delete the original test files within the repository. For instance, Figure 3 demonstrates code editing to correct a syntax error resulting from the improper use of double quotes in a Python f-string.
$\bullet$ Bash commands: Aside from the commands above, we also enable the LLM agent to directly invoke bash commands to interact with the internal environment. As shown in Figure 3, the LLM agent uses the command “export PYTHONPATH=/repo/src”.
4 Experiment
We evaluate the effectiveness of Repo2Run on 420 Python code repositories. As the popular option, we select gpt-4o-2024-05-13 for subsequent experiments, with the temperature uniformly set to 0.2.
4.1 Benchmark
To the best of our knowledge, there is no prior work similar to Repo2Run that constructs an environment for arbitrary Python repositories. To validate the capability of Repo2Run in environment configuration, we create our new benchmark consisting of selected Python repositories from GitHub based on the following criteria:
$\bullet$ Creation date: To minimize the impact of data contamination, we carefully selected repositories created in 2024, ensuring they are not part of the training data for mainstream large language models.
$\bullet$ Star count: To ensure the quality of the repositories, we only select those with a star count greater than 100.
$\bullet$ Test directory: Pytest is a leading Python testing framework that is compatible with many other Python testing tools, such as unittest. It identifies and runs test files and test functions that are named with a “test_” prefix or a “_test” suffix. Python repositories usually place their test files in the “test” or “tests” directory. To effectively filter repositories with unit tests, we keep repositories with the “test” or “tests” directory in their root directory to filter the repositories that most probably have unit tests.
Through the selection process, we crawl all 449 repositories that meet all the above requirements [^3] The crawling was performed in December 2024.. Then, we filter 420 repositories containing at least one unit test. These repositories constitute our benchmark for subsequent experiments. The details of the benchmark are presented in Appendix D.
4.2 Evaluation metrics
$\bullet$ Dockerfile Generation Success Rate (DGSR): It indicates the percentage of attempts where the method successfully generates a runnable Dockerfile. To be considered successful, the generated Dockerfile must be able to build without errors. If the Dockerfile for a code repository successfully builds, it is regarded as a successful generation. Generating runnable Dockerfile is fundamental for successfully configuring the environment.
$\bullet$ Environment Configuration Success Rate (ECSR): It represents the percentage of attempts where the method successfully configures environments. For a successful configuration, the generated Dockerfile must not only build successfully but also allow tests to run by “pytest” in the Docker container. We are only concerned with whether tests can be executed, regardless of whether they pass or fail, as outcomes of tests may inherently vary within the repository.
Table 1: The results of different baselines.
Metric DGSR ECSR
pipreqs 29.8% (125) 6.0% (25)
LLM generator 47.6% (200) 22.1% (93)
SWE-agent 26.9% (113) 9.0% (38)
Repo2Run 100% (420) 86.0% (361)
4.3 Baselines
$\bullet$ pipreqs [^4] https://github.com/bndr/pipreqs: It is an automated tool that generates a “requirements.txt” file by analyzing the import statements in the Python scripts and identifying the necessary dependencies without LLM. Using the requirements.txt file generated by pipreqs, we create a Dockerfile. The detail is provided in Appendix E.
$\bullet$ LLM generator: The “README” file in a code repository usually contains environment configuration instructions. Therefore, we directly drive the LLM to read the “README” file and generate an executable Dockerfile accordingly.
$\bullet$ SWE-agent [26]: SWE-agent establishes a custom agent-computer interface (ACI) that facilitates the LLM agent’s interaction with the repository environment by allowing actions such as reading files, editing files, and executing bash commands. Initially intended as an LLM agent for bug fixing, we preserve its framework and default settings, adjust its prompts, and use it as a baseline.
4.4 Experimental Results
The results of different baselines are presented in Table 1. We observe that Repo2Run consistently outperforms other baselines on both DGSR and ECSR. Repo2Run ultimately completed environment configuration for 361 code repositories, achieving an ECSR of 86.0%. It is 63.9% higher than the highest rate achieved by other methods, demonstrating great advantages. Due to the design of atomic configuration synthesis, Repo2Run successfully generates Dockerfiles that can be built successfully for all 420 code repositories, which other tools cannot guarantee.
For pipreqs, the main failures come from two reasons. First, generating the requirements.txt fails when there are issues within the repository, such as encoding errors or syntax errors in the files. This happens in 30 repositories (7.1%). Second, even when requirements.txt is generated, it might not download properly due to package version conflicts. This occurs in 265 repositories (63.1%). Besides, both the LLM generator and SWE-agent fail to ensure that the generated Dockerfile can be successfully built due to the lack of an ensuring mechanism. Surprisingly, the ability of SWE-agent, a general agent framework, to generate Dockerfiles is even weaker than simply letting the LLM read the “README” file. This indicates that a general agent framework cannot guarantee the generation of runnable Dockerfiles. Ensuring mechanisms like atomic configuration synthesis are necessary to effectively use the interactive information from the agent to generate runnable Dockerfiles.
Table 2: The results of ablation experiments.
Metric DGSR ECSR
w/o dual-environment 92.4% (388) 41.7% (175)
w/o Dockerfile generator 19.5% (82) 13.8% (58)
Repo2Run 100% (420) 86.0% (361)
4.5 Ablabtion of Repo2Run
To investigate the impacts of the dual-environment architecture and Dockerfile generator separately, as two parts of atomic configuration synthesis, we separately remove each component of them. For the experiment without the dual-environment architecture, we retain only the internal environment’s bash commands as the most basic interface and remove all other tools. For the experiment without the Dockerfile generator, we directly instruct the LLM to generate a runnable Dockerfile.
Experimental result of the ablation study is shown in Table 2. We observe that removing the dual-environment architecture and retaining only bash commands results in a 7.6% decrease in DGSR. The main reason for this drop is the removal of rollback and other designs, making the system more prone to entering uncertain states and subsequently failing to reproduce. In addition, ECSR shows a 44.3% decrease, primarily because the simplification of design makes it more difficult for the LLM agent to complete configuration in the internal environment. Besides, removing the Dockerfile generator directly leads to an 80.5% drop in DGSR. This indicates that having the LLM directly generate Dockerfiles is unlikely to fully follow the event history, resulting in Dockerfiles that fail to build successfully. This also directly causes a sharp decline in ECSR.
It is also observed that Repo2Run without the Dockerfile generator is outperformed by the LLM generator. This is because the LLM generator leverages the “README” file, which provides a clear, simple and high-level overview of the environment configuration, allowing for more accurate Dockerfile generation. In contrast, the event history-based approach lacks this context, making it harder for the LLM to fully understand the configuration goals. However, the Dockerfile generator effectively utilizes the detailed event history, highlighting the complementary roles of both components of Repo2Run in generating a reliable Dockerfile.
5 Discussion
5.1 Tool usage frequency
Figure 5: Tool usage frequency of the success configuration.
As shown in Figure 5, we analyze the invocation times of various action types in 361 successfully configured projects, including the five actions within the internal environment and the action of changing the base image. Bash commands are the most frequently used action, as they encompass the majority of instructions. Additionally, we observe that the LLM agent tends to call dependency installation quite frequently, averaging about 18 times per configuration, which means roughly 18 dependencies are installed per configuration on average. Moreover, the LLM agent calls test running approximately 3.5 times per configuration on average, which typically helps the agent better identify issues. Among successful configurations, we see 48 instances of changing the base image, accounting for 13.3%  of success cases, indicating that the initial selection of the base image is often incorrect and requires subsequent adjustments.
5.2 Time consumption
Figure 6: The time distribution of each successful configuration.
Figure 6 shows the distribution of time spent successfully configuring 361 code repositories. The average time for successful configuration using Repo2Run is 29.03 minutes. 111 (30.7%) of the repositories are successfully configured in less than 10 minutes. Additionally, our empirical study through sampling indicates an average manual configuration time of 21.33 minutes (Appendix A). Considering network differences and randomness, Repo2Run achieves a time consumption comparable to manual configuration. Additionally, for complex issues, Repo2Run shows greater advantages over manual configuration. Repo2Run successfully configures all the cases that were manually successful in our empirical study. Moreover, the cases where Repo2Run fails are also not successfully configured manually.
5.3 Case study
For the code repositories that fail to configure, we manually inspect the reasons for failures and find that most are due to issues within the repositories themselves. Figure 7 illustrates a “module not found” error when configuring the repository “jialuechen/deepfolio”. In this case, the issue arises from the absence of updating the unit tests. The test file “test_stats.py” attempts to import modules from “deepfolio.stats”, but “stats” does not exist in “deepfolio”. Consequently, no matter how the LLM agent operates, it cannot directly run this test. This highlights the importance for developers to continuously update existing unit tests as the code repository evolves.
6 Related Work
6.1 Environment configuration and Dockerfile generation
Oss-Fuzz-Gen [17] relies on predefined build instructions (such as “./bootstrap.sh”, “./configure”, “make”) to build projects for fuzzing but lacks of flexibility for diverse projects when specified files are absent.
Existing solutions that help developers write Dockerfiles broadly fall into three categories: (1) Template-based generators that create Dockerfiles based on project context [2, 18], (2) Task-specific tools such as DockerizeMe which supports environment dependency inference for Python projects  [11] and DockerGen  [27] for dependency recommendations based on knowledge graphs built from existing Dockerfiles, and (3) Code completion tools including GitHub Copilot [7] and HumpBack [8] that generate suggestions for developers while writing Dockerfiles. The use of deep learning models to generate Dockerfiles based on natural language specifications of software requirements is also investigated [21]. While these approaches provide valuable assistance, they either require significant manual input from developers or are limited to specific use cases.
6.2 LLM-based agent
LLM-based agents typically consist of four key components: planning, memory, perception, and action [25]. Planning is crucial for agent systems, as it schedules agents to ensure a smooth process. LLM-based agents employ various planning strategies, including single [3] or multiple planners [15], single [29] or multi-turn planning [33], and single [9] or multi-path planning [31]. The memory component in LLM-based agents stores historical data to support coherent reasoning and complex tasks. Implementations vary in terms of memory duration (short-term [6] or long-term memory [28]), memory ownership (specific [20] or shared memory [10]). For perception, LLM-based agents primarily utilize textual input [19, 28] (natural language and programming language) and visual input [24, 22] (images and diagrams) to perceive and process information. To extend capabilities beyond interactive dialogue, the action component employs various external tools [30, 16], such as searching tools, file operations, and GUI operations. Nowadays, LLM-based agents have demonstrated superior performance compared to standalone LLMs in various software engineering tasks [13, 5, 1]. However, no LLM-based agents are specifically designed for environment configuration currently. To fill this gap, this paper employs a novel approach for automated coding environment configuration and Dockerfile generation.
7 Conclusion
In this paper, we propose Repo2Run, the first LLM-based agent for automated coding environment configuration and Dockerfile generation for Python repositories based on atomic configuration synthesis. With a dual-environment architecture and a Dockerfile generator, Repo2Run is able to select and change the base image, manage and install dependencies based on action observation and rollback mechanism, and utilize bash commands and existing test suites. Our evaluation of 420 popular Python repositories hosted on GitHub demonstrates the effectiveness of Repo2Run with an 86.0% success rate.
Impact Statement
Repo2Run demonstrates the potential and impact that LLM-based agents can have towards automating coding environment configuration and docker file generation. We are hopeful that Repo2Run may inspire more works that advance broader aspects of automated software engineering. There might be more potential societal consequences of our work, none of which which we feel must be specifically highlighted here.
Appendix
Appendix A Manual experiment
A.1 Settings
@@ A.2 Survey
Regarding environment configuration, three participants indicated that they spend a significant amount of time configuring the environment when faced with an unfamiliar code repository; another three stated that, although they spend a long time, it is generally manageable; two participants reported spending minimal time.
In terms of successful environment configuration in their regular work, three participants mentioned they only fail with extremely complex environments, while five indicated they can build environments for most medium-scale repositories. As for their confidence in successfully configuring unfamiliar environments, six participants expressed that they are usually successful, and two said they are sometimes successful.
When it comes to the amount of time they are willing to wait to configure an unfamiliar repository, four participants are willing to wait for over 90 minutes, two are willing to wait 40-60 minutes, one is willing to wait 20-40 minutes, and one is only willing to wait 10-20 minutes.
Participants’ overall evaluation of configuring code running environments is diverse: one finds it very troublesome with many issues, four find it somewhat troublesome, two indicate moderate difficulty, and one finds it relatively simple.
Additionally, all participants expressed a high or very high willingness to use a tool that could automatically configure the environment for an unfamiliar repository.
A.3 Experiment guideline
@@ 1. Environment Setup
Configure the Docker environment. Verify the installation is successful: If installed correctly, you should be able to use the following command (python:3.10 is just an example; select the base image according to your requirements):
@@ 2. Overall Procedure
Our objective is to install the given package in a Docker container, configure its environment, and be able to run its internal tests. During the process, record the time developers spend configuring the environment and eventually save the logs using docker logs.
2.1 Review the Repository to be Configured (Optional)
Review the GitHub repository that needs to be configured.
@@ 2.5 Download the Repository
Select the repository to be configured and download its GitHub repository to a location (generally directly in the root directory):
@@
2.6 Configure the Environment
Now, use your skills to configure: First, enter the downloaded file directory, for example:
@@ 2.6 Configure the Environment
At this point, you can use your experience and information from the repository documentation and debugging error messages to configure. However, there are a few restrictions:
@@ 2.7 Completion and Logging
Scenario One: If “pytest --collect-only -q” runs without issues, you can then execute pytest. If pytest completes successfully, the task is done. Scenario Two: If you feel the package is extremely difficult to configure, for example exceeding your patience threshold (refer to your usual development habits), you may also terminate.
@@ 2.8 Fill Out the Form and Record Information
Waiting time Approximate value, indicating the time spent waiting for dependencies to download. Does not include time spent on decision-making and research. Provide an approximate range: ¡3 minutes 3-5 minutes 6-10 minutes 11-20 minutes 20-40 minutes 40-60 minutes 60-90 minutes 90+ minutes
Longest time-consuming process Describe the most time-consuming configuration process, such as: Downloading a specific dependency Resolving a specific error Incorrect Python version selection
Tolerance level Based on your subjective feeling and development experience. Rate the process (1-5): 1: Extremely unbearable 2: Somewhat unbearable, but manageable 3: Neutral, tolerable 4: Comfortable, no significant discomfort 5: Very comfortable, highly satisfactory configuration experience
Configuration difficulty Based on the complexity of the configuration process (ignoring time spent): Rate the process (1-5): 1: Very simple, completed with intuition and experience, no reference materials needed 2: Fairly simple, referred to basic materials (e.g., README), simple overall steps 3: Moderate difficulty, encountered some issues, but manageable 4: Difficult, required extensive debugging and configuration 5: Very difficult, needed numerous references and encountered unresolved or time-consuming issues
Materials referenced List the materials used for reference (e.g., README, internal configuration files, online searches, GPT). If none, mention “Directory”.
Biggest challenge during the process Describe the most troublesome aspect of the configuration process, such as long wait times, unclear error messages, dependency version conflicts, inability to find required software versions, etc.
@@
Biggest challenge during the process Waiting for configuration and installation
@@
Longest time-consuming process Resolving FileNotFoundError, ImportError, incorrect Python version selection
@@
A.4 Repository assignment
We randomly assigned each participant four unique code repositories that were successfully configured by Repo2Run. Additionally, each participant was assigned two code repositories that Repo2Run failed to configure. To avoid chance occurrences, each failed repository was assigned to two different participants. Below is the list of selections:
Successfully configured:

### v3

@@ B.1 Example
B.2 Safe command
Appendix C Pollution commands
Figure 10 shows two practical examples. As shown in Figure 10 (a), when we run “pip install cupy”, the command fails to execute because the CUDA environment is invalid. However, it still introduces the “fastrlock” and “numpy” packages which is not present in the original environment. In Figure 10 (b), when we run the “rm -rf /path/to/logs” command to delete the directory, some files fail to be deleted due to a lack of permission. However, the files that have already been deleted (e.g., log1.txt and log2.txt) are not restored. Such “pollution” caused by failed commands may make the environment enter the uncertain state.
Figure 10: Examples of failed command causing “pollution” to the environment.
We also identified 30 Python packages that may introduce “pollution” when installing via pip, as listed in Table LABEL:tab:pollution_example. This process can be easily reproduced in a clean Docker environment (such as the initial environment shown in Figure 10 (a)). For example, simply running “pip install adb” — even without specifying a version, which defaults to the latest — can cause such “pollution”. This demonstrates that such scenarios are quite common in real-world development and also underscores the importance of the rollback action in Repo2Run.
Table 6: 30 examples of "pollution" during pip installation. “Package_name” represents the name of the package that failed to download, "pollution" represents other packages introduced by the failed download, and "pollution" count represents the number of these introduced packages.
package_name "pollution" "pollution" count
zbarlight pillow 1
mxnet-cu91 chardet, idna, urllib3 3
texthero blis, catalogue, certifi, charset-normalizer, click, … 41
url six 1
robotframework-ride certifi, charset-normalizer, idna, numpy, packaging, … 10
adb libusb1, typing 2
onegov-core fastcache, mailthon, passlib, polib, pytz, … 12
postal six 1
changes requests 1
mxnet-cu75mkl chardet, idna, urllib3 3
winpdb numpy, six 2
slybot attrs, Automat, certifi, chardet, charset-normalizer, … 30
fbprophet aiohappeyeyeballs, aiohttp, aiosignal, appdirs, async-timeout, … 44
mxnet-cu75 chardet, idna, urllib3 3
libarchive nose 1
atari-py numpy, six 2
reppy cachetools, certifi, charset-normalizer, idna, python-dateutil, … 8
sovrin leveldb, libnacl, msgpack-python, orderedset, Pympler, … 15
scrapely numpy, six, w3lib 3
kevinsr python-version 1
pysurvive colored, numpy, pillow, psutil, pygtrie, … 6
horovod cffi, cloudpickle, packaging, psutil, PyYAML, … 6
cupy fastrlock, numpy 2
face-recognition face-recognition-models 1
nsot gunicorn, idna, ipaddress, ipython, itypes, … 15
gooey colored, numpy, pillow, psutil, pygtrie, … 6
wxpython numpy, six 2
neuralcoref annotated-types, blis, boto3, botocore, catalogue, … 47
apache-airflow-backport-providers-apache-hive apispec, argcomplete, attrs, babel, cached-property, … 58
bcolz numpy 1
Appendix D Benchmark statistic
As demonstrated in Table LABEL:tab:loc_staticstic, we list the statistics on lines of code for repositories in our constructed benchmark.
Table 7: Statistics on Lines of Code (LOC) for repositories in our benchmark. The table below illustrates the statistics regarding the lines of code (LOC) of all repositories in the benchmark. It separately counts the number of repositories with LOC greater than 10,000 and greater than 100,000. The last two rows provide reference LOC for well-known industry-level repositories.
Configuration state #LOC > 10,000 #LOC > 100,000 Medium Max
Success(361) 220 (60.9%) 27 (7.5%) 14,578 6,431,084
Fail(59) 50 (84.7%) 12 (30.6%) 43,001 3,584,542
All(420) 270 (64.3%) 39 (9.3%) 15,854 6,431,084
requests (reference) Yes No 12,782 12,782
pandas (reference) Yes Yes 571,849 571,849
Appendix E Empirical study
E.1 Manual Experiment
E.1.1 Settings
@@ E.1.2 Survey
Regarding environment building, three participants indicated that they spend a significant amount of time building the environment when faced with an unfamiliar code repository; another three stated that, although they spend a long time, it is generally manageable; two participants reported spending minimal time.
In terms of successful environment building in their regular work, three participants mentioned they only fail with extremely complex environments, while five indicated they can build environments for most medium-scale repositories. As for their confidence in successfully building unfamiliar environments, six participants expressed that they are usually successful, and two said they are sometimes successful.
When it comes to the amount of time they are willing to wait to build an unfamiliar repository, four participants are willing to wait for over 90 minutes, two are willing to wait 40-60 minutes, one is willing to wait 20-40 minutes, and one is only willing to wait 10-20 minutes.
Participants’ overall evaluation of building code running environments is diverse: one finds it very troublesome with many issues, four find it somewhat troublesome, two indicate moderate difficulty, and one finds it relatively simple.
Additionally, all participants expressed a high or very high willingness to use a tool that could automatically build the environment for an unfamiliar repository.
E.1.3 Experiment guideline
@@ 1. Environment Setup
Build the Docker environment. Verify the installation is successful: If installed correctly, you should be able to use the following command (python:3.10 is just an example; select the base image according to your requirements):
@@ 2. Overall Procedure
Our objective is to install the given package in a Docker container, build its environment, and be able to run its internal tests. During the process, record the time developers spend building the environment and eventually save the logs using docker logs.
2.1 Review the Repository to be Built (Optional)
Review the GitHub repository that needs to be built.
@@ 2.5 Download the Repository
Select the repository to be built and download its GitHub repository to a location (generally directly in the root directory):
@@
2.6 Build the Environment
Now, use your skills to build: First, enter the downloaded file directory, for example:
@@ 2.6 Build the Environment
At this point, you can use your experience and information from the repository documentation and debugging error messages to build. However, there are a few restrictions:
@@ 2.7 Completion and Logging
Scenario One: If “pytest --collect-only -q” runs without issues, you can then execute pytest. If pytest completes successfully, the task is done. Scenario Two: If you feel the package is extremely difficult to build, for example exceeding your patience threshold (refer to your usual development habits), you may also terminate.
@@ 2.8 Fill Out the Form and Record Information
Here, Table LABEL:tab:form presents the form that developers are required to complete after the environment building process, including items such as “Is it successful?”, “Final base image used”, “Reason for failure”, and so on. Besides, Tables LABEL:tab:form_example1 and LABEL:tab:form_example2 show two examples filled out by developers.
Table 8: The form that developers should fill in after the environment building.
@@
Waiting time Approximate value, indicating the time spent waiting for dependencies to download. Does not include time spent on decision-making and research. Provide an approximate range: <3 minutes 3-5 minutes 6-10 minutes 11-20 minutes 20-40 minutes 40-60 minutes 60-90 minutes 90+ minutes
Longest time-consuming process Describe the most time-consuming building process, such as: Downloading a specific dependency Resolving a specific error Incorrect Python version selection
Tolerance level Based on your subjective feeling and development experience. Rate the process (1-5): 1: Extremely unbearable 2: Somewhat unbearable, but manageable 3: Neutral, tolerable 4: Comfortable, no significant discomfort 5: Very comfortable, highly satisfactory building experience
Building difficulty Based on the complexity of the building process (ignoring time spent): Rate the process (1-5): 1: Very simple, completed with intuition and experience, no reference materials needed 2: Fairly simple, referred to basic materials (e.g., README), simple overall steps 3: Moderate difficulty, encountered some issues, but manageable 4: Difficult, required extensive debugging and building 5: Very difficult, needed numerous references and encountered unresolved or time-consuming issues
Materials referenced List the materials used for reference (e.g., README, internal building files, online searches, GPT). If none, mention “Directory".
Biggest challenge during the process Describe the most troublesome aspect of the building process, such as long wait times, unclear error messages, dependency version conflicts, inability to find required software versions, etc.
@@
Table 9: A sample form filled by a developer after environment building, illustrating a successful environment building.
@@
Biggest challenge during the process Waiting for building and installation
@@
Table 10: A sample form filled by a developer after environment building, illustrating a failed environment building.
@@
Longest time-consuming process Resolving a series of FileNotFoundError, ImportError, incorrect Python version selection
@@
E.1.4 Repository assignment
We randomly assigned each participant four unique code repositories that were successfully built by Repo2Run. Additionally, each participant was assigned two code repositories that Repo2Run failed to build. To avoid chance occurrences, each failed repository was assigned to two different participants. Below is the list of selections:
Successfully built: